Thane Company is interested in establishing the relationship between electricity costs and machine hours. Data have been collected and a regression analysis prepared using Excel. The monthly data and the regression output follow:
  

Month Machine Hours Electricity Costs
January 2,900 $ 18,600
February 3,300 $ 21,400
March 2,300 $ 13,700
April 3,500 $ 23,400
May 4,200 $ 28,450
June 3,700 $ 22,400
July 4,500 $ 24,950
August 3,900 $ 22,950
September 2,400 $ 15,900
October 4,100 $ 26,400
November 5,300 $ 31,400
December 4,900 $ 27,950

  

Summary Output
Regression Statistics
Multiple R .959
R Square .920
Adjusted R2 .912
Standard Error 1,550.32
Observations 12.00

  

Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 2,992.80 1,933.78 1.55 0.15 (1,315.92) 7,301.52
Machine Hours 5.37 0.50 10.70 0.00 4.25 6.49

  
Based on the results of the high-low analysis, the estimate of electricity costs in a month with 2,600 machine hours would be: (Round "Per unit" answer to 2 decimal places and all dollar amount to nearest whole dollar.)
Answer #1

Answer

  • First, using High Low method, we should find out:
    Variable cost per machine hour, and
    Fixed cost per month.
  • High Low working

Months

Units

Cost

High Level

Nov

                            5,300

$                   31,400.00

Low Level

Mar

                            2,300

$                   13,700.00

Difference

                            3,000

$                   17,700.00

A

Difference in Cost

$                   17,700.00

B

Difference in units

                              3,000

C = A/B

Variable cost per unit

$                           5.900

Working

High Level

Low Level

A

Total Cost

$                 31,400.00

$                   13,700.00

B

Total Units

5300

2300

C

Variable cost per unit

$                           5.90

$                             5.90

D = B x C

Total Variable cost

$                 31,270.00

$                   13,570.00

E = A - D

Total Fixed Cost

$                       130.00

$                         130.00

  • Variable cost = $ 5.90 per machine hour
    Fixed Cost = $ 130 per month.
  • Requirement:
    Electricity cost for 2600 machine hours = $ 130 fixed cost + ($5.90 x 2600 MHs)
    = 130 + 15340
    = $ 15,470 Answer